OpenStudy (anonymous):

solve the system of equation 3x=-5y 6x-7y=17

OpenStudy (anonymous):

x= 5/3; y= -1

OpenStudy (anonymous):

what did you do to get that answer

OpenStudy (anonymous):

3x=-5y (add 5y to the left side) 3x+5y=0. Then you stack the two equations on top of each other and eliminate the x variable (you can eliminate y but its easier to eliminate x). So it should look like: 3x+5y=0 6x-7y=17 Your trying to eliminate x so multiply the top equation by -2 (-2*[3x+5y=0]). It should look like this: -6x-10y=0 6x-7y=17 Adding them together you get : -17y=17 Solve for y: y=-1 Plug y into the first or second equation and solve for x 3x=-5(-1) x=5/3 Sorry it took so long for me to type. Wanted to explain it as best I could

OpenStudy (anonymous):

its okay thanks though i think you explained it better then my teacher